Day shift, 6/6

Patrick Toale (toale@cuhep.colorado.edu)
Sun, 6 Jun 1999 15:01:05 -0600 (MDT)

Shift Summary for day shift, 6/6/99

On Shift: Sydney Taegar, Pat Toale

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 6.75

Accounting of lost time: waiting on Search + Secure

========================================================================
Runs ended this shift:
Run 13392 - COSMIC
Run 13393 - LASER
Run 13394 - COSMIC

Tapes brought to FCC this shift: 0
(including tapes for online split)

========================================================================
Accesses during shift: Open access

========================================================================
Description of Shift:

7:00 - Run 13392 (COSMICS) in progress.
Hogan powers up DC2 and DC4.

7:38 - Survey crew arrives to do "as found" on DC's.

10:10 - Stopped run 13392.

10:20 - Start run 13393, LASER_SCAN.

11:05 - DC survey completed, begin to refill He bags.
Start run 13394, COSMICS.

11:30 - Magnet crash button pulled out.

13:00 - Sydney attempted a laser measurement of DC's, no instructions
could be found anywhere. Sydney's procedure and results are in the
log book, p112.

13:47 - Stopped run 13394, prepared to switch over to BEAM_MUON_4E11
configuration. Ed has changed Trig1[V*MU2] prescale from 1 to 2.

14:00 - He bags finally are full. Sydney starts closeup checklist.
Exp. Crash Switch in control room pushed in to disable
opening BEAM STOP.

14:20 - Start run 13395, BEAM_MUON_4E11 -- writing to tapes QKK001-003,
QKB001. Get 3X KUMQ DYC Busy error, fixed
by reloading Kumquats and KFREND. Also, PIPELINE error which is
fixed by reseting the pipeline. And, ETOT TW quad dropped out, ET
reloaded. SDAQ complains that ktev3a is full(99%), deleted files
in data/DC and data/CSI areas.

15:30 - Still waiting on search and secure.

N.B. Intensity should be limited to 4E11 tonight
since RickF may have to access NM2 tomorrow. MCR has
promised us 6 booster batches at this intensity. DC
crew wants to collect 15 million muons before we switch
over to hadrons.


Thanks,
Pat Toale
--------------------------------------------------------------
Graduate Research Assistant
Department of Physics University of Colorado at Boulder
Office: F-231 Phone: (303) 492-4363
E-Mail: toale@cuhep.colorado.edu